OpenStudy (anonymous):

x = 7/8 over 7/9 times 6/9 What is the value of X?

OpenStudy (akshay_budhkar):

\[(7/8)\div(7/9)\times6/9\]??

OpenStudy (akshay_budhkar):

is that your question?

OpenStudy (anonymous):

yesss thank youu

OpenStudy (akshay_budhkar):

ok.. so we have \[(7/8) \times 9/7 \times 6/9 = 6/8 = 3/4\]

OpenStudy (akshay_budhkar):

you get it?

OpenStudy (anonymous):

so x= 3/4

OpenStudy (akshay_budhkar):

yes. did you get how i solved it?

OpenStudy (anonymous):

no :( please explain it

OpenStudy (akshay_budhkar):

yes sure

OpenStudy (akshay_budhkar):

when u have \[7/8 \div 7/9 = 7/8 \times 9/7\] the fraction with which you divide , it's reciprocal is multiplied as i did above so your question transforms to \[x=7/8 \times 9/7 \times 6/9\] now just cancel the 7 and 9 you are left with 6/8 which is 3/4

OpenStudy (akshay_budhkar):

so you get it now?

OpenStudy (anonymous):

yess, ama tell you another and tell me if its right please

OpenStudy (akshay_budhkar):

yea sure my pleasure :D

OpenStudy (anonymous):

(8/9)÷(9/10)×8/10 okay wait me solve itt

OpenStudy (akshay_budhkar):

go ahead

OpenStudy (anonymous):

1/9?

OpenStudy (akshay_budhkar):

nope try again go step by step

OpenStudy (anonymous):

i dont know what to do ): because all the numbers repeat

OpenStudy (anonymous):

9/10?

OpenStudy (akshay_budhkar):

ok my bad it is 64/81 i am sleepy lol.. no it is not 9/10

OpenStudy (anonymous):

i hate math ): lol ughhhh

OpenStudy (akshay_budhkar):

math is fun :D The secret is go step by step, u will like it :D

OpenStudy (akshay_budhkar):

i need to go to sleep would u be fine if i go? can u do it? or do u need me to wait till u solve this one?

OpenStudy (anonymous):

awww, its okay thank you SO MUCH Akshay! I really appreciate it <3 goodnighttt :)

OpenStudy (akshay_budhkar):

Its my pleasure helping others :D

OpenStudy (akshay_budhkar):

when i come in next time i will see if you have solved it or not :D or else some 1 else might help you :D

OpenStudy (anonymous):

lol, yeah most probably haha <3 nightt

OpenStudy (akshay_budhkar):

night =D